### v1

Abstract
Video generation techniques have made remarkable progress, promising to be the foundation of interactive world exploration. However, existing video generation datasets are not well-suited for world exploration training as they suffer from some limitations: limited locations, short duration, static scenes, and a lack of annotations about exploration and the world. In this paper, we introduce Sekai (meaning “world” in Japanese), a high-quality first-person view worldwide video dataset with rich annotations for world exploration. It consists of over 5,000 hours of walking or drone view (FPV and UVA) videos from over 100 countries and regions across 750 cities. We develop an efficient and effective toolbox to collect, pre-process and annotate videos with location, scene, weather, crowd density, captions, and camera trajectories. Experiments demonstrate the quality of the dataset. And, we use a subset to train an interactive video world exploration model, named YUME (meaning “dream” in Japanese). We believe Sekai will benefit the area of video generation and world exploration, and motivate valuable applications.
We are looking for collaboration and self-motivated interns interested in interactive world generation. Contact: zhangkaipeng@pjlab.org.cn
44footnotetext: This work was done during the internship at Shanghai AI Laboratory.11footnotetext: Equal contribution.22footnotetext: Corresponding authors: wuyuwei@bit.edu.cn; lichuanhao@pjlab.org.cn; zhangkaipeng@pjlab.org.cn33footnotetext: Project leader.
1 Introduction
Explore. Dream. Discover.  — Mark Twain
World exploration and interaction form the foundation of humankind’s odyssey, which are practical scenarios for world generation models duan2025worldscore . These models aim to adhere to the world laws (real world or games) while facilitating unrestricted exploration and interaction within environments. In this paper, we focus on the first act of world generation—world exploration, which aims to use image, text, or video to construct a dynamic and realistic world for interactive and unrestricted exploration.
Recent advancements in video generation zhang2025packing ; yang2024cogvideox ; opensora2 ; wan2025wan ; kong2024hunyuanvideo have been remarkable, making it a promising approach for world generation through video generation. Meanwhile, camera-controlled video generation hecameractrl ; namekata2024sg ; hou2024training is a suitable way for world exploration, since camera trajectories can be converted by keyboard and mouse inputs. However, generating long and realistic videos with precise camera control remains a significant challenge. Data is a critical challenge. Existing video generation datasets nan2024openvid ; chen2024panda ; ju2024miradata are not well-suited for world exploration as they suffer from limitations: limited locations, short duration, static scenes, and a lack of annotations about exploration (e.g., camera trajectories) and world annotations (e.g., location, weather and scene).
[…]
We do dataset statistics analysis to show the scale and diversity of the dataset. And, we validate the accuracy of annotations for YouTube videos. Besides, we use a subset of the Sekai-Real to train an interactive video world exploration model, named yume ({CJK}UTF8minゆめ, meaning “dream” in Japanese).
[…]
The main contribution of this work is the dataset, Sekai. It is a high-quality long-form video dataset for worldwide exploration by walking or drone. We annotate diverse labels, including captions, location, scene, weather, crowd density, time-of-day, and camera trajectories for each video. Experiments demonstrate the effectiveness and the quality of the dataset. We believe Sekai will benefit the areas of video generation and world generation, and motivate valuable applications. We provide an introduction video on the project page.
Figure 2: An overview of the Sekai dataset. Sekai-Real is collected from YouTube with high-quality annotations, while Sekai-Game is collected from a game with ground-truth annotations.
2 Related Work
2.1 World Generation Model
Recent years have seen a growing interest in video generation zhang2025packing ; kong2024hunyuanvideo ; wang2025wan ; chen2024videocrafter2 ; li2024t2v ; xing2024dynamicrafter ; xu2024easyanimate , 3D scene generation yu2024wonderworld ; yu2024wonderjourney ; chung2023luciddreamer ; engstler2024invisible ; hollein2023text2room ; zhang2024text2nerf , and 4D generation bahmani20244d ; bahmani2024tc4d ; zhang20244diffusion ; xu2024comp4d ; zhang2025matrixgame , with significant advancements opening up new possibilities in the development of world generation models brooks2024video ; agarwal2025cosmos ; kang2024far ; meng2024towards ; xiang2024pandora . In the realm of video generation, text-to-video generation chen2024videocrafter2 ; li2024t2v has played a pivotal role, achieving high-fidelity results, while image-to-video generation xing2024dynamicrafter ; xu2024easyanimate ; shi2024motion has also seen notable advancements. Sora brooks2024video further underscores the significance of video generation in the context of world generation models. Among 3D scene generation methods, techniques chung2023luciddreamer ; engstler2024invisible ; hollein2023text2room ; yu2024wonderworld utilize depth estimation models li2024megasam ; chen2025video ; piccinelli2024unidepth to extend 2D scenes into 3D representations. 4D scene generation bahmani20244d ; bahmani2024tc4d ; zhang20244diffusion further introduces dynamics, focusing on the evolution of objects or scenes over time xu2024comp4d and dynamic interactions zhang2025matrixgame . This paper primarily focuses on interactive video generation for world exploration, aiming to construct a dynamic and realistic world using image, text, or video for unrestricted exploration.
2.2 Video Generation Dataset
The continuous development of annotated datasets has played a pivotal role in shaping the landscape of artificial intelligence-generated content, offering both insights and challenges for accurate model assessment. Existing video generation datasets can be categorized as specific-scenario and open-scenario. Typical specific-scenario datasets including UCF-101 soomro2012ucf101 , Taichi-HD siarohin2019first , SkyTimelapse xiong2018learning , FaceForensics++ rossler2019faceforensics++ , ChronoMagic yuan2024chronomagic and Celebv-HQ zhu2022celebv . These datasets have limited amount of data (with a total duration of less than 800 hours), limited individual video duration (with an average length of less than 20 seconds), and generally lack annotation information (only a few datasets, such as ChromoMagic, provide caption annotations). Open-scenario datasets bain2021frozen ; wang2023internvid ; lin2024open ; wang2020panda have somewhat alleviated issues with data scale and annotation information. For example, OpenSoraPlan-V1.0 lin2024open includes videos with a total duration of 274 hours, each accompanied by detailed captions. Similarly, the recently introduced OpenVid-1M dataset nan2024openvid comprises videos totaling 2100 hours, with long captions provided for each video. However, the average duration of individual videos still does not exceed 25 seconds, and they only provide caption annotations. MiraData ju2024miradata consists of longer videos with an average length of 72.1 seconds. It is still not long enough for the world exploration, and exploration annotations (e.g., camera poses or keyboard and mouse inputs) and world annotations (e.g., location, time and weather) are missing. By contrast, the proposed Sekai dataset focuses on egocentric world exploration, which covers walking and drone view videos across diverse locations and scenes with long video duration (1 to 39 minutes, average is 2 minutes) and rich exploratory and world annotation.
3 Dataset Curation
The overall process of curating the Sekai dataset includes four major parts: video collection, pre-processing, annotation, and sampling, seeing Figure 3 for an illustration.
Figure 3: The dataset curation pipeline. * indicates we annotate trajectories for partial data.
3.1 Video Collection
In the collection stage, we collect over 8623 hours of YouTube videos and over 40 hours of game videos from Lushfoil Photography Sim.
YouTube. We manually collect high-quality video URLs from popular YouTubers and extend them by searching additional videos using related keywords (e.g., walk, drone, HDR, and 4K). In total, we collect 10471 hours of walking videos (with stereo audio) and 628 hours of drone (FPV or UAV) videos. All videos were released over the past three years, with a 30-minute to 12-hour duration. They are at least 1080P with 30 to 60 FPS. We download the 1080P version with the highest Mbps for further video processing and annotation. Due to network issues and some videos are broken, there are 8409 hours of walking videos and 214 hours of drone videos after downloading.
Video Game. Beyond real-world data, we collect additional data from the video game since its ground-truth annotations are accessible (e.g., location, weather, and camera trajectory). Lushfoil Photography Sim is a video game that allows walking or using a first-person drone to explore real-world landscapes. It is built by Unreal Engine 5 and showcases the game’s locations in stunning visual fidelity, making it an excellent source for collecting realistic synthetic data. We use OBS Studio to record 40 hours videos at 1080P 30FPS (8 to 12 Mbps) with diverse locations and weather. Scaling the amount of data is low-cost.
3.2 Video Pre-processing
For YouTube videos, we trim two minutes from the start and end of each original video to remove the opening and ending. Then we do the following steps and obtain 6620 hours (Sekai-Real) and 60 hours (Sekai-Game) of video clips for YouTube and the game, respectively.
Shot Boundary Detection. YouTube videos are often cut and stitched, and video games commonly feature teleportation points—both of which contribute to discontinuous shot segments in one video. Thus, following Cosmos agarwal2025cosmos , we employ TransNetV2 soucek2024transnet with a threshold of 0.4 for shot boundary detection. However, the original implementation runs slowly. We refactored the codebase for GPU acceleration, which is five times faster than the original version. In particular, we use the PyNVideoCodec library for video decoding and employ the CVCUDA library to offload frame operations such as color space conversion and histogram computation to the GPU. We trim five seconds from the start and end of each shot. After shot detection, the duration of video clips is from 1 to 5.88 hours.
Clip Extraction and Transcoding. Considering practical processing, we split each shot into multiple one-minute clips (shorter than one minute will be discarded). In model training, we can stitch contiguous clips according to the computation resources. We re-encode each video clip using the PyNVideoCodec library to standardize the diverse codec configurations in the raw videos, targeting 720p at 30fps in H.265 MP4 format with a bitrate of 4 Mbps. Evaluation of the transcoded video clips across diverse scenes yields PSNR values above 35, indicating no perceptible visual degradation. We think the world exploration should contain realistic sound. Thus, we keep the audio of walking videos. We trim the audio tracks based on the timestamps of the video clips, re-encode them into AAC format using FFmpeg at 48kHz, and mux each audio clip with its corresponding video clip.
Luminance Filtering. Overly dark or bright videos are not suitable for model training. We apply a simple filter based on the luma channel in YUV color space, and remove video clips with more than 15 consecutive frames of extremely high or low average brightness. Especially, this step is necessary for video game data, as the engine often employs simplified lighting and camera systems. In this step, we filter out 300 hours of videos.
Quality Filtering. We use COVER cover2024cpvrws , a comprehensive video quality evaluator to filter low-quality video clips according to the technical quality metric. Technical quality evaluates issues such as image clarity, transmission distortion, and transcoding artifacts. The lowest-scoring 10% of video clips are removed after filtering.
Subtitle Filtering. Some videos contain hardcoded subtitles, which are artificial texts embedded in the video frames. These subtitles compromise the video’s fidelity to the real world and may introduce misleading patterns during model training. To mitigate this, we apply VideoSubFinder to detect hardcoded subtitles on the bottom one-third of the video frames. A clip is flagged if it contains any subtitle that remains visible for more than 0.75 seconds, in order to reduce false positives. All flagged clips are removed, resulting in the exclusion of approximately 5% of the video clips.
Camera Trajectory Filtering. For Sekai-Real, we employ a state-of-the-art structure from motion (SfM) model to extract camera trajectories. However, some trajectories exhibit implausible or counter-intuitive motions, so we heuristically filter out abnormal cases using the following rules. Specifically, we exclude video clips if they satisfy either of the following: (1) Multiple abrupt trajectory reversals (i.e., directional changes exceeding 150 degrees) within a 10-second window. (2) A camera viewpoint shift greater than 60 degrees between two consecutive frames. (3) A camera position displacement greater than 5 times the average displacement of the 30 consecutive frames containing these two frames. This filtering phase is only conducted for partial data annotated with camera trajectories.
3.3 Video Annotation of Sekai-Real
We annotate video clips using large vision-language models.
Location. Utilizing the Google YouTube Data API, we fetch the title and description of each video. Since most videos contain multiple chapters filmed at different locations with timeline-based descriptions, we employ GPT-4o hurst2024gpt to extract a formatted location for each chapter with the ISO 3166 country/region code attached for subsequent processing. We use the interval tree to efficiently match each video clip to its corresponding chapter based on the timestamp, thereby retrieving the location information. Video clips that cannot be uniquely matched to a chapter are discarded, which accounts for approximately 8% of the total clips.
Category and Caption. We adopt a two-stage strategy to annotate each video with category and caption. In the first stage, the video is classified along four orthogonal dimensions: scene type, weather, time of day, and crowd density, each with mutually exclusive labels. The model selects the most suitable label for each and abstains when uncertain. In the second stage, we carefully design prompts that incorporate the predicted category labels, location information, and video frames to generate detailed, time-ordered descriptions of actions and scenes for each video clip. Practically, we extract one frame every two seconds from each video clip and use 72B version of Qwen2.5-VL bai2025qwen2 to annotate them. We deploy vLLM kwon2023efficient inference services with Nginx reese2008nginx for load balance. The final caption length averages over 176 words per video clip.
Camera Trajectories. We experimented with various camera trajectory annotation methods of different types, including the visual odometry method DPVO teed2023deep , the deep visual SLAM framework MegaSaM li2024megasam , and a carefully designed 3D transformer VGGT wang2025vggt that outputs 3D quantities. Through empirical experiments and comparisons, we chose MegaSaM as the baseline annotation method and made adjustments to optimize annotation accuracy and efficiency. First, we replaced the monocular depth estimation model Depth Anything yang2024depth used in MegaSaM with Video Depth Anything chen2025video , which performs better in terms of temporal consistency. We also tried modifying the metric depth estimation model in MegaSaM by replacing UniDepth piccinelli2024unidepth with DepthPro bochkovskii2024depth . Additionally, we optimized the official implementation of MegaSaM to support cross-machine, multi-GPU parallel inference, significantly improving annotation efficiency. We annotate over 600 hours of videos (after the sampling).
3.4 Video Annotation of Sekai-Game
We developed a concise yet comprehensive toolchain based on the open-source tools RE-UE4SS and OBS Studio to capture ground-truth annotations from video games. RE-UE4SS is a powerful script system for Unreal Engine, enabling access and modification of the UE object system with minimal overhead at runtime. Based on its Lua Scripting API, we develop practical tools for video collection and annotation, including the standardization of camera system configuration, real-time camera pose capture, GUI hiding, ensuring the collection of clean data with aligned annotations.
The location and category are obtained from the description of the game map, and the prompt used for captioning is tightly modified to better suit the video game context. For camera trajectories, the captured camera poses are further calibrated to compensate for delays and interpolated to synchronize with the video frames.
3.5 Video Sampling
Given the prohibitive cost of training on the full Sekai-Real, we propose a strategy to sample the best of the best clips with the highest quality and diversity. The number is related to the computational budget for further video generation model training. In this paper, we sample 400 hours of the best of the best videos as Sekai-Real-HQ.
3.5.1 Quality Sampling
We sample the highest-quality clips according to two aspects: aesthetic quality and semantic quality. Aesthetic quality reflects the visual harmony among different elements in the video. Semantic quality assesses the semantic completeness and consistency of the content. We use COVER cover2024cpvrws to obtain two quality scores and sum them for each video clip. We sample a $\alpha_{quality}=0.7$ proportion of video clips with the highest scores.
3.5.2 Diversity Sampling
We balance the videos using the following modules one by one. And for Sekai-Real-HQ, the sampling ratio $\alpha_{content}$, $\alpha_{loc}$, $\alpha_{cate}$, $\alpha_{camera}$ are equal to $70\%$, $60\%$, $60\%$, and $75\%$, respectivelty.
Content Diversity. Given the vast volume of video clips, the presence of similar video clips is inevitable. We use InternVideo2 wang2024internvideo2 to extract embeddings for each video clip, and apply mini batch K-Means scikit-learn to cluster the embeddings of each countryregion. Subsequently, in each cluster, we use the scores in quality sampling to rank the samples. Then we iteratively sample a video clip and remove its most similar one until $1-\alpha_{content}$ proportion of video clips have been removed.
Location Diversity. We denote the number of cities as $N_{c}$. For each city, we count the number of video clips as $N$. Given a sampling ratio $\alpha_{loc}$, we sort the cities in ascending order based on their $N$. For each city in this order, we sample approximately $N\cdot\alpha_{loc}/N$ videos from each city. If it is larger than the corresponding $N$, we sample all video clips for this city and redistribute the shortfall proportionally across the remaining cities by updating $\alpha_{loc}$.
Category Diversity. To ensure broad coverage across semantic categories, we perform inverse-probability weighted sampling based on four independent categories: weather, scene, time of day, and crowd density. For each category, we compute the frequency of each label and assign sampling probabilities inversely proportional to their frequencies. Assuming independence among categories, the sampling probability for a video is initialized as the product of its label probabilities across the four categories. These probabilities are then normalized to sum to 1. We perform non-replacement sampling according to these probabilities until $\alpha_{cate}$ proportion of video clips have been sampled.
Camera Trajectory Diversity. We perform trajectory-aware sampling by the following steps. First, for the remaining videos, we calculate a direction vector (from the start to end of the trajectory) and the overall jitter, defined as the Euclidean norm of positional variance computed every 30 frames. Next, direction vectors are discretized into bins mapped onto a sphere, and jitter values are also discretized into bins. Then, a joint grouping is formed based on the direction and jitter bins. Finally, we do average sampling in each joint group according to the sampling ratio $\alpha_{camera}$.
4 Dataset Statistics
Figure 4: Statistical information on five dimensions of the Sekai-Real dataset.
For the Sekai-Game collection, we have considered the data balance issue while playing the game. Figure 4 provides an overview of the statistical information of the Sekai-Real dataset. Overall, the Sekai-Real dataset includes 101 countries and regions, with the video durations exhibiting a distinct long-tail distribution. The top eight countries (such as the Japan, the United States, and the United Kingdom) account for approximately 60% of the total duration. From various perspectives, we categorize the Sekai-Real dataset into four weather types, four scene types, four time-of-day categories, and five levels of crowd density. Specifically, the majority of the videos were outdoors. For outdoor videos, the weather conditions are predominantly sunny and cloudy, and the dataset also includes unique weather situations, such as rain and snow, increasing the diversity of the dataset. Regarding the time of filming, most videos were shot during the daytime, followed by nighttime footage. This provides a mix of well-lit scenes under natural light and darker scenes with artificial lighting, offering diverse challenges for model learning. In terms of crowd density, the distribution is relatively uniform, ranging from the sparsely populated forests of Finland to the bustling streets of Japan. This stepwise distribution covers a variety of crowd density scenarios, providing diverse training settings for the model, such as curriculum learning and evaluation of video generation models under varying levels of crowd density.
(a) Overall Video Quality Score
(b) Location Distribution
(c) Caption Tokens Distribution
(d) Camera Trajectory Jitter Distribution
Figure 5: Statistics of the proposed Sekai-Real and Sekai-Real-HQ dataset.
The statistics of Sekai-Real and Sekai-Real-HQ across multiple dimensions are shown in Figure 5. Sekai-Real-HQ is the best of the best subset of Sekai-Real, and with more balanced data distribution. Seeing Figure 5 (a), Sekai-Real demonstrates strong overall video quality scores, with more than half of the videos scoring above 0.9, while Sekai-Real-HQ exhibits a higher mean video quality score and a lower variance to address the long-tail distribution issue. Figure 5 (b) shows the distribution of Sekai-Real’s locations. Both Sekai-Real and Sekai-Real-HQ cover a wide range of countries globally. Sekai-Real-HQ demonstrates a more balanced distribution, which is more effective in mitigating potential bias during model training. Figure 5 (c) shows the distribution of caption token counts in Sekai-Real and OpenVid-1M. It can be observed that the average number of caption tokens in Sekai-Real exceeds 200, which is significantly higher than OpenVid-1M’s average of around 130. Figure 5 (d) shows the distribution of camera trajectory jitter values before and after applying Camera Trajectory Diversity. We can observe that the distribution for Sekai-Real-HQ is smoother compared to that of the Subset of Sekai-Real (500 hours, before applying Camera Trajectory Diversity). This indicates that Sekai-Real-HQ achieves better diversity and a more uniform distribution.
5 YUME Model
Figure 6: Examples of videos generated by YUME using keyboard and mouse control.
We train an interactive world exploration model named YUME ({CJK}UTF8minゆめ, meaning “dream” in Japanese) using a subset of the Sekai-Real-HQ. Specifically, it receives an image and allows unrestricted exploitation using keyboard and mouse control from users. We show some examples in Figure 6.
6 Conclusion
In this paper, we introduce a new video dataset, named Sekai, for video generation-based world exploration. It consists of over 5,000 hours of walking or drone view (FPV and UVA) videos from 101 countries and over 750 cities. We develop an efficient and effective toolchain to pre-process and annotate videos. For each video, we annotate location, scene type, weather, crowd density, captions, and camera trajectories. Besides, we introduce a video sampling module to sample the best of the best videos according to the model training budget. Experiments demonstrate the effectiveness and the quality of the dataset. We believe Sekai will benefit the area of video world generation and raise valuable applications.
7 Limitation
Insufficient training. Due to the limited computational resources, we use only a small proportion of Sekai-Real-HQ in model training.
Insufficient camera trajectory annotation. Due to our limited computational resources, for Sekai-Real, we annotate camera trajectories only for partial data.